OpenStudy (anonymous):

A(2, 1), B(2, 4) C(6, 4) and D(6, 1) are the vertices of a rectangle. a) Find the area of rectangle ABCD. b) E(3, 4) is a point on BC. Find the area of triangle AED c) What is the relationship between the area of the rectangle ABCD and the triangle AED?

OpenStudy (bahrom7893):

Just draw it out, find the length and width and multiply

OpenStudy (bahrom7893):

or u can use distance formula, but graphing is probably the best option..

OpenStudy (anonymous):

Area = AB*BC (length*width) AB = sqrt[(2-2)^2 + (4-1)^2] AB = 3 BC = sqrt[(6-2)^2 + (4-4)^2] BC = 4 area = 4*3 = 12 square units

OpenStudy (anonymous):

the best answer! i loved your step by step process! keep it up.